### v3

@@ Abstract
Detectors deployed in high-resolution neutrino experiments, particle calorimetry, or dark matter candidate searches require dense and massive active materials and, in some cases, extremely fine segmentation. This is essential for achieving precise three-dimensional tracking of the interaction products and enabling accurate particle-flow reconstruction. Organic scintillator detectors, for example, in the form of scintillating fibres, offer sub-millimetre spatial and sub-nanosecond temporal resolution. However, such systems introduce significant challenges in construction and demand a large number of readout electronics channels, leading to extremely high costs that are difficult to mitigate. In this article, we propose a paradigm shift in the detection of elementary particles that leads to ultrafast three-dimensional high-resolution imaging in large volumes of unsegmented scintillator. The key enabling technologies are plenoptic systems and time-resolving single-photon avalanche diode (SPAD) array imaging sensors.
Together, they allow us, for the first time ever with a plenoptic camera, the reconstruction of the origin of single photons in the scintillator, thereby facilitating an event-by-event analysis. A case study focused on neutrino detection demonstrates the unique potential of this approach, achieving full event reconstruction with a spatial resolution on the order of two hundred micrometres. This work paves the way for a new class of particle scintillator-based detectors, whose capabilities should be further enhanced through future developments and expanded to Cherenkov light detection and calorimetry at collider neutrino experiments, searches for neutrinoless double beta decay, as well as applications such as medical imaging and fast neutron detection.
@@
Modern elementary particle detectors deployed at high-energy physics experiments often have to cover huge volumes or surfaces and, at the same time, provide a spatial resolution on the order of 100 m with a time resolution of a few hundred picoseconds or better. Moreover, dense active materials are necessary for detecting weakly interacting particles, such as neutrinos [1, 2, 3, 4, 5, 6] and certain categories of dark matter candidates [7, 8, 9, 10], as well as for electromagnetic and hadronic calorimetry [11, 12, 13, 14, 15].
Scintillator-based detectors constitute a suitable choice for active, high-density detection systems, offering the potential for sub-millimetre spatial resolution. Organic scintillators emit blue or green isotropic light, typically in the range of 8,000 to 13,000 photons per MeV loss, when traversed by ionising radiation [16], with a decay time down to about 1-2 ns [17, 18]. Advances in detector technology have enabled 3D segmentation down to 1 cm with sub-nanosecond timing in tonne-scale scintillator volumes [19, 20, 21, 22, 23]. R&D on 3D printing of plastic scintillator is also underway [24, 25, 26, 27, 28, 29, 30].
These detectors, typically read out by Silicon Photomultipliers (SiPMs) [31], are also suitable for precise time-of-flight measurements of both charged particles [32] and neutrons [33, 34]. In the form of thin scintillating fibres, spatial resolutions below can be achieved [35, 36]. However, the main drawback lies in the challenging and costly construction, due to the very large number of channels required, particularly when targeting tonne-scale detectors.
@@
In recent years, a new class of photosensors has seen tremendous improvements: Single-Photon Avalanche Diode (SPAD) arrays. These arrays consist of multiple photosensitive diodes manufactured in CMOS technology, which are independently read out to provide single-photon images with sub-nanosecond time resolution. They are sometimes also referred to as digital SiPMs. Unlike SiPMs, SPAD arrays integrate readout electronics directly on-chip [37]. They are well-suited for applications including LIDAR [38], non-line-of-sight Imaging [39], Raman spectroscopy [40], and event cameras [41]. SPAD arrays featuring pixel pitches as small as [42, 43], or equipped with time-to-digital converters (TDCs) providing timestamp resolutions down to [38, 44], have been demonstrated in the literature. Recently, fast CMOS SPAD array sensors have been proposed as a cost-effective light readout system for scintillating fibre-based particle detectors [45].
@@
The next step in scintillator detectors involves 3D imaging of photon-starved particle events in monolithic volumes. Ref. [46] was able to reconstruct the position of gamma rays with a SPAD array sensor. However, the spatial resolution was limited by the use of the circle of confusion method, applicable only to point-like sources. Other independent works, based on Monte Carlo (MC) simulation studies, have proposed more complex optical systems adopting conventional photosensors to enhance the sensitivity to the spatial depth of particle interactions, showing potential resolutions around 1 to 10 cm, depending on the volume size: array of multiple lenses coupled to photomultiplier tubes (PMTs) for a large-scale unsegmented liquid scintillator detector [47]; coded masks [48] or objective lenses [49] coupled to SiPM to detect vacuum ultraviolet (VUV) scintillation light in liquid argon. In Ref. [50], a system of three objective lenses viewing three orthogonal faces of a few-centimetre plastic scintillator volume for fast neutron detection was studied with MC simulations.
@@
The device with the greatest potential for 3D imaging of photon-starved events is the plenoptic camera. It combines a single main objective lens (main lens) with a micro-lens array (MLA) placed in front of an imaging photosensor. Each micro-lens projected onto a subset of sensor pixels acts as a tiny camera with a slightly different perspective. This enables stereoscopic vision and thus depth reconstruction from a single exposure. A plenoptic camera is capable of capturing the “light field”, described by the function , which encodes the light intensity at each spatial point propagating in direction [51]. For this reason, it is often referred to as a light-field camera. Although the concept dates back to 1908 [52], it is only within the past two decades that the computationally intensive image post-processing has become practical [53, 54, 55]. Alternative plenoptic configurations have been proposed [56] and commercialised [57]. Recently, a plenoptic camera instrumented with a charge-coupled device (CCD) sensor has been proposed for 3D particle dosimetry [58]; however, it lacks the timing information and single-photon sensitivity required to resolve images of individual particles. Closely related to the concept of plenoptic cameras, Ref. [59, 60] investigates a design comprising lens arrays placed on the six faces of a 60 cm plastic scintillator cube, read out using SPAD array imagers, studied through MC simulations for gamma-ray detection via Compton-scattered electrons. As noted by the authors, a principal limitation of the approach is that tracks exceeding beyond the field of view of a single lens cannot be accurately reconstructed, thereby restricting the detector’s capability to the tracking of particles below a certain energy threshold. A plenoptic system is mentioned as an alternative option, but was neither conceptualised nor studied. As of now, the application of plenoptic cameras with sub-nanosecond timing to image particle events remains unexplored.
@@
In this article, we propose and demonstrate a paradigm shift in the detection of particles in a large unsegmented scintillating volume, resulting in high-resolution tracking and calorimetry of multi-particle events. The new detector consists of a system of multiple plenoptic cameras, each one with an MLA behind the main lens and a SPAD array sensor with time-resolving capabilities. A pixel pitch of the order of 10 m is desirable to constrain the light field properly. The post-processing of the 2D photon-starved images captured by the plenoptic cameras provides a pure 3D image by ray-tracing each detected photon back to its origin, leveraging the acquired light field information. If an organic scintillator with a decay time of a few nanoseconds is employed, a single-pixel timestamp resolution on the order of a few hundred picoseconds enables the effective suppression of the dominant fraction of dark counts via time-coincidence techniques. More generally, sub-nanosecond time resolution is often required in particle physics experiments to obtain a high-purity signal sample. The attenuation length, in both its scattering and absorption/emission components, shall be much longer than the size of the detector such that the original directions of the detected photons are preserved. This is easily achievable in organic scintillators with attenuation lengths of a few metres for plastics [17] and a few tens of metres for liquids [61, 62]. In this work, we demonstrate the new detector concept with the first ever-built plenoptic camera instrumented with a SPAD array, and complement it with realistic MC simulation studies. We developed novel methods for image post-processing and a data-driven calibration of the optical model. The first prototype of this detector, which we named PLATON (PLenoptic imAge of Tracked photONs), is shown in Fig. 1. Based on the results of the prototype measurement campaign, we provide the recipe of a future PLATON detector, whose development is underway, determining the design parameters of both the plenoptic system and the SPAD array sensor. We studied its design with an MC simulation of both a large cubic metre size detector as well as in a smaller PLATON module exposed to an accelerator neutrino beam. A deep neural network (NN), based on the transformer architecture developed initially for large-language models [63, 64], was developed to precisely capture both the lateral and depth spatial information with sub-millimetre resolution. This work lays the foundation for an entirely novel approach to detecting elementary particles in dense materials, offering exceptionally high spatial and temporal resolution.
@@ 2.1 The PLATON prototype
The PLATON prototype, shown in Fig. 1, adopts SwissSPAD2 [65] as the imaging sensor. It is a gated SPAD array with a 16.39 m pixel pitch, a peak photon-detection efficiency of 5% at 520 nm (photon-detection probability of 50% at 520 nm and a fill factor of 10.5%, and no on-chip per-pixel microlenses) and a median dark count rate of 0.26 counts per second (cps) per of active area. A single image frame corresponds to a gate window that can be varied between 10 ns and 100 s. The SPAD array returns single-bit frames, where an active pixel indicates that its SPAD has been activated, either by a photon or thermal noise. 2D intensity images with greyscale information are obtained from a stack of multiple frames. The plenoptic system was made in collaboration with Raytrix GmbH [57], which designed and mounted the MLA (f/2.4, micro-lens diameter of 125 m). During the assembly process, some glue leaked onto the MLA, making small areas on the left and right sides unusable for accurate imaging. This can be seen in the inset of Fig. 1. A faster MLA, such as f/1, desirable to maximise light collection efficiency, could not be used because the glob top, applied to protect the wire bonds of SwissSPAD2, was too thick to place the MLA sufficiently close to the sensor. This limitation will be overcome in a future prototype. A Voigtländer Nokton II 25 mm photographic lens, refitted with a C-Mount, was attached to the printed circuit board, on which the sensor is situated. The focus is set at a 30 cm distance to operate in a focused plenoptic camera mode, with a virtual image being formed behind the sensor. It was operated with f/2.4 to match the f-number of the MLA. With these specifications, we expect a depth resolution of the order of 1 to 5 mm and a sub-mm lateral resolution with a depth of field of 150 mm, calculated based on the effective resolution ratio (ERR) and the overall minimal change in virtual depth, as described in [56]. In the future, a better resolution will be achieved with a higher light collection efficiency and optics, currently constrained by the slower MLA f-number.
@@
The image post-processing was performed as described in Sec. 4.2. The calibration of the PLATON camera is discussed in Sec. 4.3. The results of the calibration and the spatial resolution obtained with light intensity images, as well as a function of the number of counts, are shown in Fig. 2.
@@
The reconstructed 3D position of the light source was compared to its nominal position on the stage. Justified by its size, significantly smaller than the nominal lateral resolution of the PLATON camera prototype, the source was assumed to be point-like. The spatial residual of the depth, 2D and 3D position of the light source is defined as the difference between the true (“t”) and the reconstructed position (“r”):
@@
where is the depth along the optical axis of the PLATON camera, while and are the lateral dimensions. The resolution is defined as the 68% 1-sided integral of the residual distribution. The depth, 2D lateral and 3D resolutions in the regime of fully-illuminated frames, i.e. when the light field is captured to the best capability of the camera, are, respectively, 4.2 mm, 1.8 mm and 4.5 mm for the validation sample and 2.2 mm, 0.9 mm and 2.4 mm for the calibration sample. The difference in resolution between the validation and calibration data was found to be affected by remounting the camera on the movement stage between the calibration and validation data runs. This can be seen in the shift in the lateral reconstruction in Fig. 2. Thus, we will refer to the calibration sample to highlight the intrinsic resolution of the PLATON prototype.
@@
The spatial resolution was also studied as a function of the number of detected photons. A data-driven procedure was adopted: 1,000 photon-starved images were obtained by sampling the pixel position of a certain predetermined number of counts, using the intensity image of the point light source at a specific position as a PDF. With 10 detected photons, the 2D lateral resolution falls below . As expected, the 3D resolution is dominated by the depth one and becomes worse than 10 mm for fewer than 10 photons. The obtained spatial resolution is consistent with what is expected from the design specifications of the PLATON camera, including the reduced FoV (see Sec. 2.1), as well as with results obtained with the optical simulation (see Sec. 4.1). Using the simulated equivalent of the calibration sample, the extracted depth resolution is 2.25 mm, with a 2D radial resolution of 0.5 mm. When accounting for possible systematic uncertainties inherent in the stage and camera setup, these values are considered consistent with those obtained from measurements.
@@
An alternative method for the position reconstruction of the point light source was also tested on the same data samples, augmented with additional data points (see section 4 for details). We refer to it as “likelihood mehod”. It is based on the computation of the binary cross-entropy () using the data-driven noise-subtracted PDFs for different 3D positions in the movement stage. A more detailed description of the analysis can be found in Sec. 4.4. A depth resolution of 1.5 mm was achieved along with a lateral resolution of 0.1 mm. With the sampled photon-starved images, the depth and lateral resolutions decreased to, respectively, below 10 mm and 0.3 mm for 5 photons. For fewer than 10 photons, depth ambiguities arise between points located along a line extending from the principal point, as these produce nearly concentric image patterns. It is worth noting that the likelihood study should not be interpreted as the achieved spatial resolution, as it would be impractical to build a PDF for every single possible point down to the relevant precision in a relatively large volume. On the other hand, the largely improved resolution compared to the reference method described in Sec. 2.2 (better than a factor of 2 or 3 for more than 30 counts) indicates that the 2D frames provided by the PLATON prototype contain more spatial information than what can be exploited with the developed post-processing method. Such a feature highlights that a more advanced post-processing, such as one that uses prior information of the optical model more detailed than that described in Sec. 4.2 or deep learning (see Sec. 2.3.2 and 4.5) would further improve the 3D spatial resolution. Our near-future plan is to pursue both directions further.

### v4

@@ Abstract
Neutrino detectors, particle calorimeters and some dark matter detectors require dense and massive active materials. An extremely fine segmentation is desirable to achieve precise three-dimensional particle tracking. However, such systems introduce significant challenges in construction and demand a large number of readout electronics channels, leading to extremely high costs. In this article, we propose an alternative approach to elementary particle detection that enables ultrafast three-dimensional high-resolution imaging in large volumes of unsegmented scintillator. Enabling technologies are plenoptic systems and time-resolving single-photon avalanche diode array imaging sensors. Together, they enabled us, using a plenoptic camera, to reconstruct the origin of single photons in the scintillator. A case study focused on neutrino detection demonstrates full event reconstruction with a spatial resolution of two hundred micrometres. This work paves the way for a class of particle detectors whose capabilities should be further enhanced through future developments and expanded to Cherenkov light detection, medical imaging and neutron detection.
@@
Modern elementary particle detectors deployed at high-energy physics experiments often have to cover huge volumes or surfaces and, at the same time, provide a spatial resolution on the order of 100 m with a time resolution of a few hundred picoseconds or better. Moreover, dense active materials are necessary for detecting weakly interacting particles, such as neutrinos [Hyper-Kamiokande:2025fci, Hyper-Kamiokande:2018ofw, Abi:2020qib, Abe:2019vii, NOvA:2022see, MINOS:2008hdf] and certain categories of dark matter candidates [Zani:2024ybb, DarkSide-20k:2024yfq, Aprile:2020vtw, Aalbers:2022dzr], as well as for electromagnetic and hadronic calorimetry [calice, CALICE:2018ibt, 2137105, CMSHCAL:2007zcq, CMS:2009cmd].
Scintillator-based detectors constitute a suitable choice for active, high-density detection systems, offering the potential for sub-millimetre spatial resolution. Organic scintillators emit blue or green isotropic light, typically in the range of 8,000 to 13,000 photons per MeV loss, when traversed by ionising radiation [Kolanoski:2020ksk], with a decay time down to about 1-2 ns [eljen-catalogue-2025, Artikov:2005mg]. Advances in detector technology have enabled 3D segmentation down to 1 cm with sub-nanosecond timing in tonne-scale scintillator volumes with wavelength shifting (WLS) fibre readout [Sgalaberna:2017khy, sfgd-testbeam-cern, Mineev:2018ekk, Alekseev:2022jki, Boyarintsev_2021]. R&D on 3D printing of plastic scintillator is also underway [son2018improved, KIM20202910, Berns:2020ehg, Berns_2022, Sibilieva:2022ket, KAPLON2022106864, Weber:2023joy]. Another notable scintillator-based technology, currently under development, aims to get rid of the geometrical segmentation by deploying organic scintillating materials with a short optical scattering length to stochastically contain the visible photons within small localised volumes [LiquidO:2019mxd]. These detectors, typically read out by Silicon Photomultipliers (SiPMs) [hamamatsu:mppc], are also suitable for precise time-of-flight measurements of both charged particles [Korzenev:2021mny] and neutrons [Agarwal:2022kiv, Gwon:2022bix]. In the form of thin scintillating fibres, spatial resolutions below can be achieved [Joram:2015ymp, Papa:2023uqv]. However, the main drawback lies in the challenging and costly construction, due to the very large number of channels required, particularly when targeting tonne-scale detectors. However, the main drawback lies in the challenging and costly construction due to the very large number of analog readout channels required, each needing to be instrumented with corresponding digitisation electronics, particularly when targeting tonne-scale detectors.
@@
In recent years, a new class of photosensors has seen tremendous improvements: Single-Photon Avalanche Diode (SPAD) arrays. These arrays consist of multiple photosensitive diodes manufactured in CMOS technology, which are independently read out to provide single-photon images with sub-nanosecond time resolution. They are sometimes also referred to as digital SiPMs. Unlike SiPMs, SPAD arrays integrate readout electronics directly on-chip [Rochas_2003_First_fully_integrated]. Consequently, a single data line allows the readout of millions of pixels, rather than requiring an independent analogue channel and digitisation chain for each. They are well-suited for applications including LIDAR [DBLP:journals/jssc/ZhangLAPWC19], non-line-of-sight Imaging [spad-nonlineofsightimaging], Raman spectroscopy [spad-biophotonics], and event cameras [10655318]. SPAD arrays featuring pixel pitches as small as [Morimoto_OPEX:20, Shimada:22], or equipped with time-to-digital converters (TDCs) providing timestamp resolutions down to [DBLP:journals/jssc/ZhangLAPWC19, Fischer:2022sfe], have been demonstrated in the literature. Recently, fast CMOS SPAD array sensors have been proposed as a cost-effective light readout system for scintillating fibre-based particle detectors [Franks:2023ttv].
@@
The next step in scintillator detectors involves 3D imaging of photon-starved particle events in monolithic volumes. Ref. [Bocchieri_2024_Scintillation_event_imaging] was able to reconstruct the position of gamma rays with a SPAD array sensor. However, the spatial resolution was limited by the use of the circle of confusion method, applicable only to point-like sources. Other independent works, based on Monte Carlo (MC) simulation studies, have proposed more complex optical systems adopting conventional photosensors to enhance the sensitivity to the spatial depth of particle interactions, showing potential resolutions around 1 to 10 cm, depending on the volume size: array of multiple lenses coupled to photomultiplier tubes (PMTs) for a large-scale unsegmented liquid scintillator detector [Dalmasson:2017pow]; coded masks [infn-coded-mask] or objective lenses [CICERO2025170801] coupled to SiPM to detect vacuum ultraviolet (VUV) scintillation light in liquid argon. In Ref. [Pisanti_2024], a system of three objective lenses viewing three orthogonal faces of a few-centimetre plastic scintillator volume for fast neutron detection was studied with MC simulations.
@@
The device with the greatest potential for 3D imaging of photon-starved events is the plenoptic camera. It combines a single main objective lens (main lens) with a micro-lens array (MLA) placed in front of an imaging photosensor. Each micro-lens projected onto a subset of sensor pixels acts as a tiny camera with a slightly different perspective. This enables stereoscopic vision and thus depth reconstruction from a single exposure. A plenoptic camera is capable of capturing the “light field”, described by the function , which encodes the light intensity at each spatial point propagating in direction [Levoy_1996_Light_field_rendering]. For this reason, it is often referred to as a light-field camera. Although the concept dates back to 1908 [Lippmann_1908_Epreuves_reversibles_donnant], it is only within the past two decades that the computationally intensive image post-processing has become practical [Adelson_1992_Single_lens_stereo, 2005_Ng_Handheld_lightfield, 2006_Ng_Digital_Lightfield_Photography]. Alternative plenoptic configurations have been proposed [Perwass_2012_Single_lens_3D_camera] and commercialised [raytrix]. Recently, a plenoptic camera instrumented with a charge-coupled device (CCD) sensor has been proposed for 3D particle dosimetry [Goulet_2014_Novel_full_3D]; however, it lacks the timing information and single-photon sensitivity required to resolve images of individual particles. Closely related to the concept of plenoptic cameras, Ref. [10.1117/12.2055951, Lebrun:2015gH] investigates a design comprising lens arrays placed on the six faces of a 60 cm plastic scintillator cube, read out using SPAD array imagers, studied through MC simulations for gamma-ray detection via Compton-scattered electrons. As noted by the authors, a principal limitation of the approach is that tracks exceeding beyond the field of view of a single lens cannot be accurately reconstructed, thereby restricting the detector’s capability to the tracking of particles below a certain energy threshold. A plenoptic system is mentioned as an alternative option, but was neither conceptualised nor studied. As of now, the application of plenoptic cameras with sub-nanosecond timing to image particle events remains unexplored.
@@
In this article, we propose and demonstrate a paradigm shift in the detection of particles in a large unsegmented scintillating volume, resulting in high-resolution tracking and calorimetry of multi-particle events. The detector consists of a system of multiple plenoptic cameras, each one with an MLA behind the main lens and a SPAD array sensor with time-resolving capabilities. A pixel pitch of the order of 10 m is desirable to constrain the light field properly. The post-processing of the 2D photon-starved images captured by the plenoptic cameras provides a pure 3D image by ray-tracing each detected photon back to its origin, leveraging the acquired light field information. If an organic scintillator with a decay time of a few nanoseconds is employed, a single-pixel timestamp resolution on the order of a few hundred picoseconds enables the effective suppression of the dominant fraction of dark counts via time-coincidence techniques. More generally, sub-nanosecond time resolution is often required in particle physics experiments to obtain a high-purity signal sample. The attenuation length, in both its scattering and absorption/emission components, shall be much longer than the size of the detector such that the original directions of the detected photons are preserved. This is easily achievable in organic scintillators with attenuation lengths of a few metres for plastics [eljen-catalogue-2025] and a few tens of metres for liquids [Djurcic:2015vqa, ABUSLEME2021164823]. In this work, we demonstrate a detector concept with a plenoptic camera instrumented with a SPAD array, and complement it with realistic MC simulation studies. We developed different methods for image post-processing and a data-driven calibration of the optical model. The first prototype of this detector concept, which we named PLATON-prototype (PLenoptic imAge of Tracked photONs), is shown in Fig. 1. Based on the results of the prototype measurement campaign, we provide the recipe of a future PLATON detector, whose development is underway, determining the design parameters of both the plenoptic system and the SPAD array sensor. We studied its design with a MC simulation of both a large cubic metre size detector as well as in a smaller PLATON-10cm module exposed to an accelerator neutrino beam. A deep neural network (NN), based on the transformer architecture developed initially for large-language models [vaswani2017attention, 10278387], was developed to precisely capture both the lateral and depth spatial information with sub-millimetre resolution. This work lays the foundation for an entirely different approach to detecting elementary particles in dense materials, offering exceptionally high spatial and temporal resolution.
@@ 2.1 The PLATON prototype
The PLATON-prototype, shown in Fig. 1, adopts SwissSPAD2 [Ulku_2019_A_512_x] as the imaging sensor. It is a gated SPAD array with a 16.39 m pixel pitch, a peak photon-detection efficiency of 5% at 520 nm (photon-detection probability of 50% at 520 nm and a fill factor of 10.5%, and no on-chip per-pixel micro-lenses) and a median dark count rate of 0.26 counts per second (cps) per of active area. A single image frame corresponds to a gate window that can be varied between 10 ns and 100 s. The SPAD array returns single-bit frames, where an active pixel indicates that its SPAD has been activated, either by a photon or thermal noise. 2D intensity images with greyscale information are obtained from a stack of multiple frames. The plenoptic system was made in collaboration with Raytrix GmbH [raytrix], which designed and mounted the MLA (f/2.4, micro-lens diameter of 125 m). During the assembly process, some glue leaked onto the MLA, making small areas on the left and right sides unusable for accurate imaging. This can be seen in the inset of Fig. 1. A faster MLA, such as f/1, desirable to maximise light collection efficiency, could not be used because the glob top, applied to protect the wire bonds of SwissSPAD2, was too thick to place the MLA sufficiently close to the sensor. This limitation will be overcome in a future prototype. A Voigtländer Nokton II 25 mm photographic lens, refitted with a C-Mount, was attached to the printed circuit board, on which the sensor is situated. The focus is set at a 30 cm distance to operate in a focused plenoptic camera mode, with a virtual image being formed behind the sensor. It was operated with f/2.4 to match the f-number of the MLA. With these specifications, we expect a depth resolution of the order of 1 to 5 mm and a sub-mm lateral resolution with a depth of field of 150 mm, calculated based on the effective resolution ratio (ERR) and the overall minimal change in virtual depth, as described in [Perwass_2012_Single_lens_3D_camera]. In the future, a better resolution will be achieved with a higher light collection efficiency and optics, currently constrained by the slower MLA f-number.
@@
The image post-processing was performed as described in Sec. 4.2. The calibration of the PLATON-prototype is discussed in Sec. 4.3. The results of the calibration and the spatial resolution obtained with light intensity images, as well as a function of the number of counts, are shown in Fig. 2.
@@
The reconstructed 3D position of the light source was compared to its nominal position on the stage. Justified by its size, significantly smaller than the nominal lateral resolution of the PLATON-prototype, the source was assumed to be point-like. The spatial residual of the depth, 2D and 3D position of the light source is defined as the difference between the true (“t”) and the reconstructed position (“r”):
@@
where is the depth along the optical axis of the PLATON-prototype, while and are the lateral dimensions. The resolution is defined as the 68% 1-sided integral of the residual distribution. The depth, 2D lateral and 3D resolutions in the regime of fully-illuminated frames, i.e. when the light field is captured to the best capability of the camera, are, respectively, 4.2 mm, 1.8 mm and 4.5 mm for the validation sample and 2.2 mm, 0.9 mm and 2.4 mm for the calibration sample. The difference in resolution between the validation and calibration data was found to be affected by remounting the camera on the movement stage between the calibration and validation data runs. This can be seen in the shift in the lateral reconstruction in Fig. 2. Thus, we will refer to the calibration sample to highlight the intrinsic resolution of the PLATON-prototype.
@@
The spatial resolution was also studied as a function of the number of detected photons. A data-driven procedure was adopted: 1,000 photon-starved images were obtained by sampling the pixel position of a certain predetermined number of counts, using the intensity image of the point light source at a specific position as a PDF. With 10 detected photons, the 2D lateral resolution falls below . As expected, the 3D resolution is dominated by the depth one and becomes worse than 10 mm for fewer than 10 photons. The obtained spatial resolution is consistent with what is expected from the design specifications of the PLATON-prototype, including the reduced FoV (see Sec. 2.1), as well as with results obtained with the optical simulation (see Sec. 4.1). Using the simulated equivalent of the calibration sample, the extracted depth resolution is 2.25 mm, with a 2D radial resolution of 0.5 mm. When accounting for possible systematic uncertainties inherent in the stage and camera setup, these values are considered consistent with those obtained from measurements.
@@
An alternative reconstruction method for point-like sources was also tested using a similarly acquired data sample. This method is based on evaluating the likelihood of the measured light patterns with respect to a reference data sample. A depth resolution of 1.5 mm was achieved along with a lateral resolution of 0.1 mm. With the sampled photon-starved images, the depth and lateral resolutions decreased to, respectively, below 10 mm and 0.3 mm for 5 photons. For fewer than 10 photons, depth ambiguities arise between points located along a line extending from the principal point, as these produce nearly concentric image patterns. For further details, see Sec. 4.4. It is worth noting that the likelihood study should not be interpreted as the achieved spatial resolution, as it would be impractical to build a PDF for every single possible point down to the relevant precision in a relatively large volume. On the other hand, the largely improved resolution compared to the reference method described in Sec. 2.2 (better than a factor of 2 or 3 for more than 30 counts) indicates that the 2D frames provided by the PLATON-prototype contain more spatial information than what can be exploited with the developed post-processing method. Such a feature highlights that a more advanced post-processing, such as one that uses prior information of the optical model more detailed than that described in Sec. 4.2 or deep learning (see Sec. 2.3.2 and 4.5) would further improve the 3D spatial resolution. Our near-future plan is to pursue both directions further.